### v1

# Deception Against Data-Driven Linear-Quadratic Control

Abstract
Deception is a common defense mechanism against adversaries with an information disadvantage. It can force such adversaries to select suboptimal policies for a defender’s benefit. We consider a setting where an adversary tries to learn the optimal linear-quadratic attack against a system, the dynamics of which it does not know. On the other end, a defender who knows its dynamics exploits its information advantage and injects a deceptive input into the system to mislead the adversary. The defender’s aim is to then strategically design this deceptive input: it should force the adversary to learn, as closely as possible, a pre-selected attack that is different from the optimal one. We show that this deception design problem boils down to the solution of a coupled algebraic Riccati and a Lyapunov equation which, however, are challenging to tackle analytically. Nevertheless, we use a block successive over-relaxation algorithm to extract their solution numerically and prove the algorithm’s convergence under certain conditions. We perform simulations on a benchmark aircraft, where we showcase how the proposed algorithm can mislead adversaries into learning attacks that are less performance-degrading.
{IEEEkeywords}
Deception, data-driven control, linear-quadratic control, successive over-relaxation.
1 Introduction
Deception is a mechanism defenders often use in adversarial environments, particularly as a tool to mislead their opponents into adopting oblivious policies. For example, adversarial entities can employ deception in cyber-physical settings as a means to render their attacks unnoticed [1, 2], but also, in other contexts such as robotics [3], autonomous vehicles [4], and warfare [5]. Despite the heterogeneity of these applications, a common theme among them is that of information asymmetry; that is, to be able to deceive the opponent, one must have an information advantage against them, and the opponent should usually be oblivious to this advantage.
In this paper, we consider deception in the context of data-driven linear-quadratic control. This is a recently emerged field in control theory that aims to design optimal controllers for systems whose dynamics are either uncertain or completely unknown. Its underlying idea is to use state and/or input histories of the system at hand to synthesize the desired control policy, instead of solving model-based equations that require strong prior empirical knowledge [6, 7]. Yet, a major backdoor in data-driven control is its sensitivity to perturbations often present in the measured data. Such perturbations can be benign when they conform to specific random patterns or are sufficiently small [8, 9]. However, this is not the case when they are specifically crafted by an opponent with an information advantage as a means for deception. Crafting such deception is the focus of this paper.
More specifically, we consider a setting where an adversary wants to compute the optimal linear quadratic attack against a dynamical system. However, being an outsider, the adversary has no prior knowledge regarding the system’s dynamics. It thus needs to either perform a direct data-driven method to learn the optimal attack or employ system identification and then learn the optimal attack indirectly. On the other end, the defender of the system is aware of the adversary’s malicious objective and is motivated to perform deception to mislead them towards learning a more benign attack. To that end, the defender exploits its information advantage with respect to the system’s dynamics and injects deceptive feedback into the system to distort the data the adversary is using. A key characteristic of such feedback is that it has the same effect on the attack that the adversary will ultimately learn, irrespective of the adversary’s actual learning algorithm.
To obtain the defender’s deceptive input, we cast its objective as a constrained optimization problem balancing two objectives: i) the desire to steer the adversary’s learned policy towards an a priori chosen benign gain; and ii) the desire to use as small of deceptive feedback as possible to maintain the nominal closed-loop performance and stability. We mathematically specify the conditions under which the defender’s optimization is well-defined and characterize its solution as the root of a coupled algebraic Riccati equation and a Lyapunov equation. Subsequently, since these coupled equations are difficult to analytically solve, we tackle them numerically with a successive over-relaxation iterative algorithm. In addition, we prove this algorithm’s convergence under certain conditions. We finally note that the proposed deception design also applies to the easier problem of deceiving minimizing data-driven linear-quadratic regulators.
Related work
Despite data-driven control being a relatively young field, relevant studies regarding its deception have recently appeared in the literature. For example, related to the philosophy of this paper is the concept of deception in data-driven Nash equilibrium seeking [10]. It considers a setting where an agent engaged in a multi-player game has an information advantage and uses this advantage to design a deceptive input that steers the other agents’ policies toward a deceptive equilibrium. Nevertheless, the setting considered in [10] is tailored to repeated static games.
The concept of deceiving data-driven methods is also often related to poisoning such methods, owing to the data corruption that implicitly or explicitly takes place as the means for deception. This idea of poisoning has attracted much interest in the machine learning community [11, 12, 13, 14, 15] and a few studies have also investigated it from a systems and control theoretic perspective. For instance, relevant examples include poisoning virtual reference feedback controllers [16, 17] and data predictive control [18]. Nevertheless, most of these techniques revolve around distantly related setups with discrete-time systems, different objectives, and sometimes without theoretical convergence guarantees. Relevant research initiatives have also focused on increasing resilience against poisoning attacks [19, 20, 21, 22]. However, the problem of explicitly deceiving a learning-based optimal control algorithm, so that it converges to a prespecified policy, has not been studied.
Moving target defense is another mechanism used to deal with adversaries [23, 24, 25]. Its main idea is that to prevent an adversary from stealthily attacking a dynamical system, it is in the interest of the defender to constantly alter the system’s dynamics to reveal or hinder the attack. However, moving target defense does not exploit the defender’s information advantage to explicitly deceive the adversary’s policy. This is the case in the deception method we develop, where one can prescribe how the defender’s deceptive input will affect the learned adversarial policy and hence steer it towards an a priori chosen benign target. While worst-case min-max approaches as in [26, 27, 28] aim to protect against fully-informed adversaries, we exploit the defender’s advantage to strategically control the adversary’s policy.
A preliminary version of this paper appeared in [29], which studied the problem of deception against minimizing data-driven linear quadratic regulators. However, [29] only provided simulation results to support its algorithm’s ability in extracting the optimal deception policy. On the other hand, hereon, we adopt a stricter approach and theoretically prove the algorithm’s convergence properties. In addition, we shed light on the problem of deceiving maximizing regulators. This is a much more challenging problem owing to the associated value function possibly becoming unbounded and thus undefined [26].
2 Preliminaries
Notation: We denote as $\mathbb{R}$ the set of real numbers, and as $\mathbb{N}$ the set of natural numbers including zero. The vector $e_{i}\in\mathbb{R}^{n}$ denotes a unit vector with a unity located in the $i$-th row. For any matrix $X$, $\left\lVert X\right\rVert_{\mathrm{F}}$ denotes its Frobenius norm, whereas $\mathrm{Ran}(X)$ denotes its range. In addition, $[X]_{ij}$ denotes the entry in its $i$-th row and $j$-th column. Moreover, we follow the matrix differentiation convention $\left[\frac{\mathrm{d}}{\mathrm{d}X}\right]_{ij}=\frac{\mathrm{d}}{\mathrm{d}[X]_{ij}}$. For a square matrix $X$, $\alpha(X)$ denotes its spectral abscissa, and $\textrm{tr}(X)$ its trace. For a symmetric matrix $X$, $\lambda_{\mathrm{min}}(X)$ and $\lambda_{\mathrm{max}}(X)$ denote its minimum and maximum eigenvalue, respectively. We denote $X\succ 0$ (respectively, $X\succeq 0)$ if $X$ is positive definite (respectively, positive semidefinite). We denote as $\textrm{vec}(X)$ the vectorization of $X$. We use $I_{n}$ to denote the identity matrix of order $n$ (we omit the subscript when the order is obvious). Finally, we denote the Kronecker delta function as $\delta_{i,j}$, so that $\delta_{i,j}=1$ if $i=j$, and $\delta_{i,j}=0$ else.
Definition 1.
Consider the following algebraic Riccati equation:
$A^{\mathrm{T}}P+PA+Q-P(B_{1}R_{1}^{-1}B_{1}^{\mathrm{T}}-B_{2}R_{2}^{-1}B_{2}^{\mathrm{T}})P=0,$ (1)
where $A,P,Q\in\mathbb{R}^{n\times n}$, $R_{1}\in\mathbb{R}^{m_{u}\times m_{u}}$, $R_{2}\in\mathbb{R}^{m_{a}\times m_{a}}$, $B_{1}\in\mathbb{R}^{n\times m_{u}}$, $B_{2}\in\mathbb{R}^{n\times m_{a}}$, $R_{1},R_{2},Q\succ 0$, and $(A,B_{1})$ is stabilizable. When it exists, we denote the unique stabilizing solution of (1) as $P\in\mathbb{S}_{+}$, with $\mathbb{S}_{+}$ being the set of positive-definite stabilizing solutions. Such a solution is positive definite, though not necessarily uniquely, but is the minimal such solution of (1) [26].
3 Problem Formulation
3.1 Closed-Loop System
Consider the continuous-time, closed-loop system for all $t\geq 0$:
$\dot{x}(t)=Ax(t),$ (2)
where $x(t)\in\mathbb{R}^{n}$ is the state of the system, and $A\in\mathbb{R}^{n\times n}$ is the closed-loop state matrix. By closed loop, we mean that $A=A_{o}+B_{u}F$, where $A_{o}\in\mathbb{R}^{n\times n}$ is the open-loop state matrix, $B_{u}\in\mathbb{R}^{n\times m_{u}}$ is the input matrix, and $F\in\mathbb{R}^{m_{u}\times n}$ is a stabilizing gain such that $A$ is Hurwitz. For example, $F$ could be a gain obtained from a pole-placement procedure that moves the eigenvalues of $A$ to the open left-half plane. For such a gain to be feasible, $(A,B_{u})$ must also be controllable.
3.2 The Adversary’s Model and Objectives
We assume that an adversary can launch an adversarial attack on system (2). In the presence of such an attack, system (2) would take the form
$\dot{x}(t)=Ax(t)+B_{a}a(t),$ (3)
with $B_{a}\in\mathbb{R}^{n\times m_{a}}$ being the adversary’s input matrix, and $a(t)\in\mathbb{R}^{m_{a}}$ being the attack. The adversary’s purpose is to then design the attack to balance two objectives: i) disturbing the stabilization of (3); and ii) moderating the attack’s magnitude. We capture this dual objective with the maximizing linear-quadratic control problem:
$\max_{K\in\mathbb{R}^{m_{a}\times n}}~{}$ $\int_{0}^{\infty}(x^{\mathrm{T}}(\tau)Qx(\tau)-a^{\mathrm{T}}(\tau)Ra(\tau))\mathrm{d}\tau,$
s.t. $\dot{x}(t)=Ax(t)+B_{a}a(t),$ (4)
$a(t)=Kx(t),$
where $Q,R\succ 0$ are weighting matrices and $K$ is the adversary’s attack gain that is to be optimized. Provided that $\lambda_{\textrm{min}}(R)$ is above a threshold, it then follows from standard linear systems theory [26] that the solution to this problem is given by
$K^{\star}=R^{-1}B_{a}^{\mathrm{T}}P,$ (5)
where $P\in\mathbb{S}_{+}$ is the minimal positive definite solution of the algebraic Riccati equation (ARE)
$A^{\mathrm{T}}P+PA+Q+PB_{a}R^{-1}B_{a}^{\mathrm{T}}P=0.$ (6)
We now present the information structure of the adversary.
Assumption 1.
The following hold true.
  1. 1.
    The adversary does not know the values of the entries of the system’s matrices $A,B_{a}$.
  2. 2.
    The adversary can measure the state of the system.
The main implication of the adversary’s information structure is that it impedes them from directly solving the ARE (6), and hence from obtaining the optimal attack gain $K^{\star}$ in (5). In that regard, the adversary is forced to gather measurements from the system over a given window and use them to either solve (4) in a direct data-driven manner, or to perform system identification and then solve (4).
3.3 The Defender’s Model and Objectives
We assume that the defender of the system is aware that a potential adversary might be gathering information from the system with the aim of synthesizing the optimal attack gain $K^{\star}$ in (5). In that respect, the defender opts to spoof the system’s dynamics during this window, and deceive the adversary into learning a prespecified gain different from $K^{\star}$. Specifically, during this window, the system’s dynamics take the form
$\dot{x}(t)=Ax(t)+B_{u}u(t)+B_{a}a(t),$ (7)
where $u(t)\in\mathbb{R}^{m_{u}}$ is the defender’s deceiving input added to spoof the system dynamics and deceive the adversary, forcing them to learn a gain $\bar{K}\in\mathbb{R}^{m_{a}\times n}$ that is different from $K^{\star}$.
We now present the information structure of the defender.
Assumption 2.
The following hold true.
  1. 1.
    The defender knows the system matrices $A,B_{u},B_{a}$.
  2. 2.
    The defender can measure the state of the system.
  3. 3.
    The defender does not know the learning algorithm the adversary is using, but knows its objective.
The fact that the defender does not know the learning algorithm of the adversary, though realistic, makes it difficult to gauge how the deceiving input $u(t)$ will affect where that algorithm will converge. An exception is if the defender chooses $u$ as
$u(t)=\Lambda x(t),$
where $\Lambda\in\mathbb{R}^{m_{u}\times n}$ is an appropriately chosen deception gain. Such a strategy simply redefines the system’s internal dynamics, and hence will generate the same effect to the learned adversarial gain irrespective of the underlying learning dynamics. Hence, if the adversary employs a learning-based scheme to estimate the optimal gain $K^{\star}$ that solves (4) while under the proposed deceptive measure, the adversary will instead learn the gain that solves
$\max_{K\in\mathbb{R}^{m_{a}\times n}}~{}$ $\int_{0}^{\infty}(x^{\mathrm{T}}(\tau)Qx(\tau)-a^{\mathrm{T}}(\tau)Ra(\tau))\mathrm{d}\tau,$
s.t. $\dot{x}(t)=(A+B_{u}\Lambda)x(t)+B_{a}a(t),$ (8)
$a(t)=Kx(t),$
where we note that the trajectories in this maximization problem are over the spoofed dynamics, whereas the trajectories in (4) were over the dynamics of the original system. Following again standard linear systems theory, one can show that the optimal solution to (8) is
$K_{u}(\Lambda)=R^{-1}B_{a}^{\mathrm{T}}P_{u}(\Lambda),$ (9)
where $P_{u}(\Lambda)\in\mathbb{R}^{n\times n}\in\mathbb{S}_{+}$ is now the minimal positive definite solution of the “spoofed” ARE:
$(A+B_{u}\Lambda)^{\mathrm{T}}P_{u}+P_{u}(A+B_{u}\Lambda)+Q\\ +P_{u}B_{a}R^{-1}B_{a}^{\mathrm{T}}P_{u}=0.$ (10)
Figure 1 depicts this deceptive process.
[…]
Figure 1: The deception scheme. The adversary uses state data $x(t)$ and input data $a(t)$ to learn the optimal attack (5) against the closed-loop system (3). However, during the adversary’s data-gathering process, the defender feeds into the closed loop a deceptive gain $\Lambda$, forcing the adversary to learn the incorrect attack (9) instead.
The above equations imply that the defender can predict the adversarial gain $K_{u}(\Lambda)$ that the enemy will ultimately learn, particularly by exploiting its information advantage with respect to the system’s dynamics. In addition, the defender can perform this prediction without using any data. Nevertheless, choosing the deception gain $\Lambda$ to control $K_{u}(\Lambda)$ is far from being a trivial task, because the dependence of $K_{u}(\Lambda)$ on $\Lambda$ goes through the minimal solution of the ARE (10). In that regard, if the defender wants, for example, to suppress the adversarial attack by forcing $K_{u}(\Lambda)$ to be very close to zero, then it is not clear how it should select $\Lambda$ to achieve this suppression optimally. Motivated by this issue, we formalize the problem we investigate in the remainder of the paper as follows.
Problem 1.
Design $\Lambda$, so that the gain $K_{u}(\Lambda)$ that the adversary will learn closely resembles a different gain $\bar{K}\in\mathbb{R}^{m_{a}\times n}$ that is prechosen by the defender.
Remark 1.
Another variation of the deception problem considered in this paper is one where a defender tries to learn the minimizing linear-quadratic regulator for the system, and an adversary tries to deceive them towards converging to an unrelated, suboptimal gain. We note that this problem is easier owing to its inherently convex nature, and one can solve it using tools identical to those presented in the following sections; see Section 6.3 for more details.
Remark 2.
While we assume the defender has full state feedback, the approach can also be extended to the output feedback case through the use of an observer or estimator. However, estimation errors might affect the impact of deception.
4 Deception Design
4.1 Deception Design using a Distance-Based Metric
One method the defender could use to design its deception gain $\Lambda$, is so that $K_{u}(\Lambda)$ is as close as possible—in the matrix norm sense—to $\bar{K}$. In other words, the defender could select its gain $\Lambda$ so that it minimizes $\left\lVert K_{u}(\Lambda)-\bar{K}\right\rVert_{\mathrm{F}}^{2}$. Under this objective, the optimal deception gain would be derived as the solution of the constrained optimization
$\inf_{\Lambda\in\mathbb{R}^{m_{u}\times n},P_{u}\in\mathbb{R}^{n\times n}}J(\Lambda,P_{u})=\left\lVert K_{u}(\Lambda)-\bar{K}\right\rVert_{\mathrm{F}}^{2}+\mathrm{tr}(\Lambda^{\mathrm{T}}\Gamma\Lambda),$
[…]
$\qquad\qquad\quad~{}P_{u}\in\mathbb{S}_{+},$
where $\bar{K}$ is as in Problem 1, $\Gamma\succ 0$, and $\mathrm{tr}(\Lambda^{\mathrm{T}}\Gamma\Lambda)$ is a regulation term. This regulation term is necessary to ensure that the minimum exists, but also to penalize how large the deception gain $\Lambda$ can be to minimize performance loss and maintain stability [^1] This is for the case where $\bar{K}$ is not stabilizing. For the case where $\bar{K}=0$, regularization is not needed; see Section 6.1. . As an alternative, the distance of the target gain $\bar{K}$ from the original gain $K^{\star}$ should be sufficiently small, assuming $K^{\star}$ exists. We provide an explicit – though conservative – condition on the distance $\left\lVert K^{\star}-\bar{K}\right\rVert_{\mathrm{F}}$ as well as the regulation weight $\Gamma$, for ensuring closed-loop stability and the feasibility of (11), in the following Lemma. The Lemma’s proof, as well as the proofs of all technical results, are postponed to the Appendix.
Lemma 1.
Let $S\in\mathbb{R}^{n\times n}$ be the unique positive-definite solution of
$(A+B_{a}K^{\star})^{\mathrm{T}}S+S(A+B_{a}K^{\star})+2I=0.$ (12)
Choose $\Gamma$ and the target gain $\bar{K}$ so that
$\left\lVert K^{\star}-\bar{K}\right\rVert_{\mathrm{F}}<\frac{\sqrt{\lambda_{\textrm{min}}(\Gamma)}}{\sqrt{\lambda_{\textrm{min}}(\Gamma)}\left\lVert SB_{a}\right\rVert_{\mathrm{F}}+\left\lVert SB_{u}\right\rVert_{\mathrm{F}}}.$ (13)
Then, (11) admits a global minimizer $(\Lambda^{\star},P_{u}^{\star})$, i.e., its infimum is a minimum. In addition, $A+B_{u}\Lambda^{\star}+B_{a}K_{u}(\Lambda^{\star})$ is strictly stable.
We note that the condition of Lemma 1, though sufficient, is not necessary for maintaining stability and feasibility. It can often prove to be quite conservative, and merely having a sufficiently large regulation matrix $\Gamma$ in terms of eigenvalues should suffice both in practice and theory. Moreover, simulation results and an additional analysis in Section 6.2 show that even this condition is often redundant, as for certain reasonable choices of $\bar{K}$ the regulation matrix $\Gamma$ can be arbitrarily small; see Section 6.1 for further details.
To proceed, we use the Karush–Kuhn–Tucker conditions to characterize the solution to the optimization problem (11), which yields the optimal deception policy for the defender.
Theorem 1.
Let $(\Lambda^{\star},P_{u}^{\star})\in\mathbb{R}^{m_{u}\times n}\times\mathbb{R}^{n\times n}$ be the optimal solution to (11). Then, $P_{u}^{\star}\in\mathbb{S}_{+}$, and there exists $\Pi^{\star}\in\mathbb{R}^{n\times n}$ symmetric such that $(\Lambda,P_{u},\Pi)=(\Lambda^{\star},P_{u}^{\star},\Pi^{\star})$ solve the system of equations
$(A+B_{u}\Lambda)^{\mathrm{T}}P_{u}+P_{u}(A+B_{u}\Lambda)$
$\qquad\qquad\qquad\qquad+Q+P_{u}B_{a}R^{-1}B_{a}^{\mathrm{T}}P_{u}=0,$ (14a)
$(A+B_{u}\Lambda+B_{a}R^{-1}B_{a}^{\mathrm{T}}P_{u})\Pi$
$+\Pi(A+B_{u}\Lambda{+}B_{a}R^{-1}B_{a}^{\mathrm{T}}P_{u})^{\mathrm{T}}-\bar{K}^{\mathrm{T}}R^{-1}B_{a}^{\mathrm{T}}$
$-B_{a}R^{-1}\bar{K}+P_{u}B_{a}R^{-2}B_{a}^{\mathrm{T}}+B_{a}R^{-2}B_{a}^{\mathrm{T}}P_{u}=0,$ (14b)
$\Lambda=-\Gamma^{-1}B_{u}^{\mathrm{T}}P_{u}\Pi.$ (14c)
According to Theorem 1, we can extract the optimal solution to the deception design problem (11) by solving the matrix equations (14a)-(14c). However, a close inspection of these equations reveals that they are not easy to solve analytically, even in the scalar case. This is mainly due to their coupled and nonlinear nature, arising from the cross-term $P_{u}B_{u}\Lambda$ in (14a), and the cross-terms $P_{u}\Pi$ and $\Lambda\Pi$ in (14b).
A potential solution to the aforementioned bottleneck could be to employ off-the-shelf numerical solvers for nonlinear systems of equations on (14a)-(14c), combined with standard heuristics that may aid their chances of success. However, not only would such a choice be computationally expensive since the number of unknowns scales quadratically with $n$ in (14a)-(14c), but it would also have no theoretical guarantees of success. A more thoughtful numerical approach is thus clearly required.
4.2 Proposed Numerical Solution
To be able to numerically attack equations (14a)-(14c) in an efficient manner, it is important to first understand their structure. Even though as a whole this system of equations does not seem to fall within any known class of systems, each of its equations – when viewed independently – has a relatively familiar structure. More specifically, equation (14a) is an ARE when viewed with $P_{u}$ being its sole unknown variable, and such AREs are straightforward to solve for their minimal positive definite solution. In addition, equation (14b) is a Lyapunov equation (LE) when viewed with $\Pi$ being its sole unknown variable, and hence has a unique solution – as long as $P_{u}$ is minimal – that we can extract even analytically. Finally, provided that both $P_{u}$ and $\Pi$ are known a priori, (14c) is a very simple computation that yields the deception gain $\Lambda$.
Algorithm 1 Block Successive Over-relaxation Algorithm to solve (14a)-(14c)
Input: Tolerance $\epsilon>0$, step size $\omega\in(0,2)$, regulation weight $\Gamma\succ 0$, desired gain $\bar{K}$, parameters $A,B_{u},B_{a},R,Q$.
      Output: Estimated solution $(\hat{\Lambda},\hat{P}_{u},\hat{\Pi})$ to (14a)-(14c).
1:procedure
2:     Initialize $\Lambda^{0}$ so that $J(\Lambda^{0},P_{u}(\Lambda^{0}))\leq J(0,P_{u}(0))$.
3:     $i\leftarrow 0$.
4:     while $i=0$ or $\left\lVert\Lambda^{i}-\Lambda^{i-1}\right\rVert_{\mathrm{F}}\geq\epsilon\cdot\omega$ do
5:         Solve for $P_{u}^{\textrm{GS}}\in\mathbb{S}_{+}$ from the ARE
$(A+B_{u}\Lambda^{i})^{\mathrm{T}}P_{u}^{\textrm{GS}}+P_{u}^{\textrm{GS}}(A+B_{u}\Lambda^{i})\\ +Q+P_{u}^{\textrm{GS}}B_{a}R^{-1}B_{a}^{\mathrm{T}}P_{u}^{\textrm{GS}}=0.$ (15)
6:         Set $P_{u}^{i+1}\leftarrow P_{u}^{\textrm{GS}}$.
7:         Solve for $\Pi^{\textrm{GS}}$ from the LE
$(A+B_{u}\Lambda^{i}+B_{a}R^{-1}B_{a}^{\mathrm{T}}P_{u}^{i+1})\Pi^{\textrm{GS}}$
${+}\Pi^{\textrm{GS}}(A{+}B_{u}\Lambda^{i}{+}B_{a}R^{-1}B_{a}^{\mathrm{T}}P_{u}^{i+1})^{\mathrm{T}}{-}\bar{K}^{\mathrm{T}}R^{-1}B_{a}^{\mathrm{T}}$ (16)
$-B_{a}R^{-1}\bar{K}+P_{u}^{i+1}B_{a}R^{-2}B_{a}^{\mathrm{T}}+B_{a}R^{-2}B_{a}^{\mathrm{T}}P_{u}^{i+1}=0.$
8:         Set $\Pi^{i+1}\leftarrow\Pi^{\textrm{GS}}$.
9:         Compute $\Lambda^{\textrm{GS}}$ as
$\Lambda^{\textrm{GS}}=-\Gamma^{-1}B_{u}^{\mathrm{T}}P_{u}^{i+1}\Pi^{i+1}.$ (17)
10:         Set $\Lambda^{i+1}\leftarrow\Lambda^{i}+\omega(\Lambda^{\textrm{GS}}-\Lambda^{i})$.
11:         $i\leftarrow i+1$.
12:     end while
13:     $\hat{\Lambda}\leftarrow\Lambda^{i}$, $\hat{P}_{u}\leftarrow P_{u}^{i}$, $\hat{\Pi}\leftarrow\Pi^{i}$.
14:end procedure
In view of the above, a potential structured numerical method to solve (14a)-(14c) is the following: one can proceed to begin with an arbitrary gain $\Lambda$, then solve (14a) for its minimal positive definite solution $P_{u}$ (provided it exists), then solve (14b) for its symmetric solution (provided it exists), then update $\Lambda$ according to (14c), and then repeat this process iteratively until a fixed point is found. Such a fixed point will necessarily be a solution to (14a)-(14c), and hence qualify for being a stationary point of (11).
The aforementioned method is known as the Gauss-Seidel method in numerical linear algebra [30]. We present a variation of it in Algorithm 1, known as block successive over-relaxation, wherein the update of the gain $\Lambda$ at each iteration goes through a low-pass filter, and where the iterates (15)-(17) imitate the structure of equations (14a)-(14c). Unlike the Gauss-Seidel and the Jacobi method (Chapter 7 in [30]), which may diverge depending on the problem’s parameters, successive over-relaxation has provable convergence guarantees as we show in Section 5.
Remark 3.
As further detailed in Section 3, the execution of Algorithm 1 does not require any data. The reason is that the defender has knowledge of the system’s dynamics, and can use these to directly predict how the deceptive gain will affect the policy the attacker will learn.
5 Convergence Analysis
We first presented Algorithm 1 in the preliminary study [29], which designed poisoning actuation attacks against defenders learning the minimizing linear-quadratic control of an unknown system. While simulation results in [29] showcased that Algorithm 1 performs well in a variety of settings, it is difficult to use such empirical results to argue about convergence properties, mainly because fixed-point iteration schemes are notorious for often being divergent. For this reason, here we will adopt a theoretically strict approach, and mathematically prove that Algorithm 1 leads to a stationary point of the deception problem (11). In addition, we will show that this stationary point, under certain conditions, corresponds to a minimum.
Towards this end, notice that in the optimization problem (11), the ARE constraint (10) completely dictates the choice of the matrix $P_{u}$. That is, once we select a specific gain $\Lambda$, then we subsequently obtain the matrix $P_{u}$ by solving the equality constraint (10) for its minimal positive definite solution. Taking this into consideration, it is possible to define the cost function of (11) uniquely with respect to $\Lambda$ as
$\tilde{J}(\Lambda)=J(\Lambda,P_{u}(\Lambda)),$ (18)
where, with a slight abuse of notation, $P_{u}(\cdot)$ is now a function with unknown analytical form, which maps $\Lambda$ to the minimal positive definite solution of the ARE
$(A+B_{u}\Lambda)^{\mathrm{T}}P_{u}(\Lambda)+P_{u}(\Lambda)(A+B_{u}\Lambda)\\ +Q+P_{u}(\Lambda)B_{a}R^{-1}B_{a}^{\mathrm{T}}P_{u}(\Lambda)=0.$ (19)
In light of this definition, in what follows we prove the following central result: one can interpret the fixed point iteration presented in Algorithm 1 as a gradient iteration on the cost function $\tilde{J}(\Lambda)$, scaled by $\Gamma^{-1}$.
Lemma 2.
For any $\Lambda\in\mathbb{R}^{m_{u}\times n}$ for which the ARE (19) admits a solution $P_{u}(\Lambda)\in\mathbb{S}_{+}$, we have
$\frac{\mathrm{d}\tilde{J}(\Lambda)}{\mathrm{d}\Lambda}=2\left(\Gamma\Lambda+B_{u}^{\mathrm{T}}P_{u}(\Lambda)\Pi(\Lambda)\right),$ (20)
where $\Pi(\Lambda)$ is a matrix function of $\Lambda$, and the unique symmetric solution of the LE
$(A+B_{u}\Lambda+B_{a}R^{-1}B_{a}^{\mathrm{T}}P_{u}(\Lambda))\Pi(\Lambda)$
$+\Pi(\Lambda)(A+B_{u}\Lambda{+}B_{a}R^{-1}B_{a}^{\mathrm{T}}P_{u}(\Lambda))^{\mathrm{T}}{-}\bar{K}^{\mathrm{T}}R^{-1}B_{a}^{\mathrm{T}}$ (21)
$-B_{a}R^{-1}\bar{K}+P_{u}(\Lambda)B_{a}R^{-2}B_{a}^{\mathrm{T}}+B_{a}R^{-2}B_{a}^{\mathrm{T}}P_{u}(\Lambda)=0.$
According to the notation and results of Lemma 2, from lines 9-10 of Algorithm 1, we have
$\omega(\Lambda^{\mathrm{GS}}-\Lambda^{i})$ $=\omega(-\Gamma^{-1}B_{u}^{\mathrm{T}}P_{u}^{i+1}\Pi^{i+1}-\Lambda^{i})$
$=\omega(-\Gamma^{-1}B_{u}^{\mathrm{T}}P_{u}(\Lambda^{i})\Pi(\Lambda^{i})-\Lambda^{i})$
$=-\frac{1}{2}\omega\Gamma^{-1}\cdot 2(B_{u}^{\mathrm{T}}P_{u}(\Lambda^{i})\Pi(\Lambda^{i})+\Gamma\Lambda^{i})$
$=-\frac{1}{2}\omega\Gamma^{-1}\frac{\mathrm{d}\tilde{J}(\Lambda^{i})}{\mathrm{d}\Lambda^{i}}.$
Hence, we can interpret the iterative update of the deception gain $\Lambda^{i}$ in line 10 of Algorithm 1 as
$\Lambda^{i+1}\leftarrow\Lambda^{i}-\frac{1}{2}\omega\Gamma^{-1}\frac{\mathrm{d}\tilde{J}(\Lambda^{i})}{\mathrm{d}\Lambda^{i}},$ (22)
where we recall that $\tilde{J}(\Lambda)$ is the cost function in (11) when $P_{u}$ is viewed implicitly as a function of $\Lambda$ (see (18)). This means that Algorithm 1 is a first-order method on the “generalized” cost (18) and should, in principle, converge to a stationary point, which would then correspond to a stationary point of problem (11). However, a couple of issues must be considered.
The first issue is that the domain of $\tilde{J}$ is not equal to all of $\mathbb{R}^{m_{u}\times n}$. This is because $\Lambda$ must be such that a solution $P_{u}(\Lambda)\in\mathbb{S}_{+}$ for (19) exists, and this is clearly not the case for all $\Lambda$ in $\mathbb{R}^{m_{u}\times n}$. Accordingly, if Algorithm 1 happens to abruptly exit the domain of $\tilde{J}$ at some iteration, it will terminate prematurely with an error while attempting to solve (15). The second issue is that even if the iterates $\Lambda^{i}$ do not violate the domain of $\tilde{J}$ throughout the execution of Algorithm 1, convergence to a stationary point can be assured only as long as the gradient of $\tilde{J}$ has a bounded Lipschitz constant [31]. Fortunately, in what follows, we will show that neither of those two issues will become an obstacle to convergence, under certain conditions.
Towards this end, let us define the set
$\mathcal{S}_{0}=\{\Lambda\in\mathbb{R}^{m_{u}\times n}~{}|~{}\exists P_{u}(\Lambda)\in\mathbb{S}_{+}~{}\textrm{and}~{}\tilde{J}(\Lambda)\leq\tilde{J}(0)\},$ (23)
where we observe that $\Lambda^{0}$, i.e., the starting point of Algorithm 1, lies inside $\mathcal{S}_{0}$. The following auxiliary lemma shows that both $P_{u}(\Lambda)$ and $\Pi(\Lambda)$ are bounded uniformly in $\mathcal{S}_{0}$, and that $A+B_{u}\Lambda+B_{a}K_{u}(\Lambda)$ remains strictly stable uniformly in $\mathcal{S}_{0}$.
Lemma 3.
Let the following strengthened version of (13) hold:
$\left\lVert K^{\star}-\bar{K}\right\rVert_{\mathrm{F}}\leq\frac{\epsilon}{2}\frac{\sqrt{\lambda_{\textrm{min}}(\Gamma)}}{\sqrt{\lambda_{\textrm{min}}(\Gamma)}\left\lVert SB_{a}\right\rVert_{\mathrm{F}}+\left\lVert SB_{u}\right\rVert_{\mathrm{F}}},$ (24)
where $S$ is as in (12), and $\epsilon\in(0,1)$. Then, the following hold.
  1. 1.
    $A+B_{u}\Lambda+B_{a}K_{u}(\Lambda)$ is strictly stable uniformly in $\mathcal{S}_{0}$. Specifically, $\sup_{\Lambda\in\mathcal{S}_{0}}\alpha(A+B_{u}\Lambda+B_{a}K_{u}(\Lambda))<0$, and, more strongly,
    $(A+B_{u}\Lambda+B_{a}K_{u}(\Lambda))^{\mathrm{T}}S\\ +S(A+B_{u}\Lambda+B_{a}K_{u}(\Lambda))\preceq-2(1-\epsilon)I$ (25)
    for all $\Lambda\in\mathcal{S}_{0}$, with $S$ as in (12).
  2. 2.
    There exist finite constants $\bar{P},~{}\bar{\Pi}>0$ such that $\left\lVert P_{u}(\Lambda)\right\rVert_{\mathrm{F}}\leq\bar{P}$ and $\left\lVert\Pi(\Lambda)\right\rVert_{\mathrm{F}}\leq\bar{\Pi}$ for all $\Lambda\in\mathcal{S}_{0}$.
The practical interpretation of Lemma 3 is that, by continuity, $P_{u}(\Lambda)\in\mathbb{S}_{+}$ exists not only everywhere in $\mathcal{S}_{0}$ but also in a neighborhood around it. Therefore, the negated gradient $-\frac{\mathrm{d}\tilde{J}(\Lambda)}{\mathrm{d}\Lambda}$ points in a direction towards which $P_{u}(\Lambda)\in\mathbb{S}_{+}$ exists, and hence this direction is in the interior of $\mathcal{S}_{0}$. Accordingly, Algorithm 1 will always remain within a set where both equations (15) and (16) admit unique solutions and will not terminate prematurely. This effectively deals with one of the two issues raised in the preceding paragraphs.
Next, we focus on the second issue raised in the preceding discussion; that the gradient of $\tilde{J}$ must admit a global Lipschitz constant on $\mathcal{S}_{0}$ for Algorithm 1 to converge to a stationary point of (11). Generally speaking, such a conclusion cannot be directly obtained by inspection because $P_{u}(\Lambda)$ and $\Pi(\Lambda)$ in (20) are derived as solutions of nonlinear equations, and hence can depend sublinearly on $\Lambda$. Nevertheless, using the uniform boundedness result asserted in the second part of Lemma 3, we are able to establish such property as follows.
Lemma 4.
Let (24) hold. Then, the gradient of the cost function $\tilde{J}(\Lambda)$, i.e., $\frac{\mathrm{d}\tilde{J}(\Lambda)}{\mathrm{d}\Lambda}$, is globally Lipschitz on $\Lambda\in\mathcal{S}_{0}$.
Finally, we combine the previous Lemmas to obtain the main result of this subsection, which characterizes the convergence properties of Algorithm 1. In particular, we show that Algorithm 1 converges to a stationary point of (11), and that this point is a minimum under a stricter condition on the regulation matrix $\Gamma$.
Theorem 2.
Let (24) hold and $\omega<\frac{2}{L}\frac{\lambda_{\mathrm{min}}(\Gamma^{-1})}{\lambda_{\mathrm{max}}^{2}(\Gamma^{-1})}$, where $L$ is the Lipschitz constant of $\frac{\mathrm{d}\tilde{J}(\Lambda)}{\mathrm{d}\Lambda}$ on $\mathcal{S}_{0}$. Consider the sequence of matrices $\{P_{u}^{i},\Pi^{i},\Lambda^{i}\}_{i\in\mathbb{N}}$ generated by Algorithm 1. Then:
  1. 1.
    At each step $i\in\mathbb{N}$, there exist unique solutions $P_{a}^{\mathrm{GS}}\in\mathbb{S}_{+},~{}\Pi^{\mathrm{GS}},~{}\Lambda^{\mathrm{GS}}$ to the equations (15)-(17);
  2. 2.
    $\tilde{J}(\Lambda^{i+1})\leq\tilde{J}(\Lambda^{i})$ for all $i\in\mathbb{N}$;
  3. 3.
    $\lim_{i\rightarrow\infty}\frac{\mathrm{d}\tilde{J}(\Lambda^{i})}{\mathrm{d}\Lambda}=0$, that is, the stationarity equations (14a)-(14c) hold in the limit, and Algorithm 1 terminates for any tolerance $\epsilon>0$;
  4. 4.
    There exists $\gamma^{\star}>0$, such that if $\Gamma\succ\gamma^{\star}I$ then Algorithm 1 converges to a minimum of (11).
Remark 4.
An alternative to having a small constant step size $\omega$ in Algorithm 1, would be to instead have an iteration-varying step $\omega^{i}$ with the property that $\sum_{i\in\mathbb{N}}(\omega^{i})^{2}<\infty$ and $\sum_{i\in\mathbb{N}}\omega^{i}=\infty$. While such a choice retains convergence guarantees, empirical results indicate that it makes Algorithm $1$ converge substantially more slowly.
6 Discussion
In this section, we discuss how we can extend the results of the preceding sections to tackle more general cases of deception in linear-quadratic control.
6.1 On the Condition of Lemma 1
A sufficient condition for guaranteeing the existence of a minimizer to (11) is inequality (13), requiring that the attack gain $\bar{K}$ the defender wants to induce be sufficiently close to the original gain $K^{\star}$. As stated in the preceding sections, such a condition can be conservative because we derived it without making any a priori assumptions regarding the deceptive target gain $\bar{K}$. On the other hand, if we consider the possibility that $\bar{K}$ will be better for the defender in terms of closed-loop performance when compared to the original gain $K^{\star}$, then (11) is unnecessary. We illustrate this in the following example, wherein the solution to the ARE (6) that yields the nominal attack $K^{\star}$ does not even exist, because $\lambda_{\textrm{min}}(R)$ is small and hence $K^{\star}$ is destabilizing.
Example 1.
Consider a scalar system where the closed-loop state matrix is $A=-1$, the input matrices are $B_{u}=B_{a}=1$, and the state cost matrix is $Q=1$. Then, then ARE (6) reduces to the scalar quadratic equation
$P^{2}-2RP+R=0.$ (26)
When $R>1$ solution, the solutions to this equation are given by $P_{1,2}=R\pm\sqrt{R-1}\sqrt{R}$, and hence the minimal positive definite solution is $P=R-\sqrt{R-1}\sqrt{R}$. Therefore, from (5), the optimal attack gain is $K^{\star}=1-\sqrt{1-R^{-1}}$. On the other hand, when $10<R<$ 1, (26) admits no solutions, the value of (4) is unbounded, and the optimal attack $K^{\star}$ destabilizing. Yet, the defender can deceive the attacker into a milder attack that renders the closed-loop stable, while bypassing the rationale of Lemma 1.
More specifically, let $R=0.5<1$ and suppose that the target attack gain the defender wants to induce is $\bar{K}=0.2$; a milder target than $K^{\star}$ which does not disrupt closed-loop stability. Then, assuming $\Gamma=0$ and thus further violating (13), the defender would need to find the gain $\Lambda$ such that $K_{u}(\Lambda)=R^{-1}P_{u}(\Lambda)=\bar{K}=0.2$, and thus $P_{u}(\Lambda)=0.1$. Plugging this in (10) yields $\Lambda=-4.1$. Therefore, $A+B_{u}\Lambda+B_{a}K_{u}(\Lambda)=A+B_{u}\Lambda+B_{a}\bar{K}=-4.9$, i.e., the minimum of (11) is attained and (10) admits a stabilizing solution. In addition, $A+B_{a}K_{u}(\Lambda)=-1+0.2=-0.8$, i.e., the deceptive attack gain does not disrupt closed-loop stability.
The example above showcases that even when (13) does not hold, (11) and its ARE constraint (10) can still be well-defined. The main idea is that it is not that the target $\bar{K}$ needs to be close to $K^{\star}$ per se, as (13) requires; rather, that it needs to be equal to an attack gain that is milder than $K^{\star}$, or at least not much worse than $K^{\star}$. Accordingly, in these cases, one should select the initialization $\Lambda^{0}$ in Algorithm $1$ to render $A+B_{u}\Lambda^{0}$ sufficiently stable, i.e., with a sufficiently small spectral absissca [^2] This implicitly requires that $(A,B_{u})$ be controllable.. In this way, one can guarantee that the first ARE iterate (15) will admit a well-defined stabilizing solution, and that (16) will admit a unique solution, too.
We capture the intuition that (13) does not need to hold when $\bar{K}$ is milder than $K^{\star}$ in the technical result below, for the case $\bar{K}=0$.
Proposition 1.
Let $\bar{K}=0$ and suppose $\mathrm{Ran}(B_{a})\subseteq\mathrm{Ran}(B_{u})$. Then, there exists $\gamma^{\star}>0$ such that if $\Gamma\prec\gamma^{\star}I$ then (11) admits a global minimizer.
Remark 5.
The condition $\mathrm{Ran}(B_{a})\subseteq\mathrm{Ran}(B_{u})$ requires that the adversary does not have more channels to attack the system than the defender has to control it.
6.2 On the Assumption of Known $Q$ and $R$ Matrices
In Assumption 2, we assumed the defender does not know the learning algorithm used by the adversary, but knows its objective, i.e., its matrices $Q$ and $R$. A natural question thus regards how well the deception scheme will work when the actual adversary matrices are different.
To investigate this question, suppose that the actual weighing matrices in the adversary’s objective are equal to $\hat{Q}\neq Q$ and $\hat{R}\neq R$. In addition, consider the cost function (11), but under these actual weighing matrices $\hat{Q},~{}\hat{R}$ of the adversary:
$\hat{J}(\Lambda)=\left\lVert\hat{K}_{u}(\Lambda)-\bar{K}\right\rVert_{\mathrm{F}}^{2}+\mathrm{tr}(\Lambda^{\mathrm{T}}\Gamma\Lambda),$ (27)
where $\hat{K}_{u}(\Lambda)=\hat{R}^{-1}B_{a}^{\mathrm{T}}\hat{P}_{u}(\Lambda)$ and
$(A+B_{u}\Lambda)^{\mathrm{T}}\hat{P}_{u}(\Lambda)+\hat{P}_{u}(\Lambda)(A+B_{u}\Lambda)+\hat{Q}\\ +\hat{P}_{u}(\Lambda)B_{a}\hat{R}^{-1}B_{a}^{\mathrm{T}}\hat{P}_{u}(\Lambda)=0,~{}\hat{P}_{u}(\Lambda)\in\mathbb{S}_{+}.$ (28)
In the following proposition, we show that when $\bar{K}=0$, the gain the defender derives by optimizing (11) is still efficient in deceiving the adversary towards $\bar{K}=0$, despite the incorrectly assumed matrices $Q$ and $R$ in (11). In particular, we show that i) in the case the actual matrices of the adversary satisfy $\hat{Q}\prec Q$ and $\hat{R}\succ R$, then the optimal value of (11) provides an upper bound for (27); and ii) in the general case, the optimal value of (11) provides an approximate upper bound for (27).
Proposition 2.
Let $\bar{K}=0$. Denote as $\Lambda^{\star}$ the minimizer of (11). The following hold true.
  1. 1.
    If $\hat{Q}\preceq Q$ and $\hat{R}\succeq R$ then $\hat{J}(\Lambda^{\star})<\tilde{J}(\Lambda^{\star})$.
  2. 2.
    Suppose $\mathrm{Ran}(B_{a})\subseteq\mathrm{Ran}(B_{u})$. Then, for every $\epsilon>0$ there exists $\gamma^{\star}>0$ such that if $\Gamma\preceq\gamma^{\star}I$ then $\hat{J}(\Lambda^{\star})\leq\tilde{J}(\Lambda^{\star})+\epsilon$.
6.3 On the Deception against Minimizing Data-Driven Linear-Quadratic Regulators
A dual setup in data-driven linear quadratic control is one where a defender wants to learn the solution to the minimization problem
$\min_{N\in\mathbb{R}^{m_{u}\times n}}~{}$ $\int_{0}^{\infty}(x^{\mathrm{T}}(\tau)Qx(\tau)+u^{\mathrm{T}}(\tau)Mu(\tau))\mathrm{d}\tau,$
s.t. $\dot{x}(t)=Ax(t)+B_{u}u(t),$ (29)
$u(t)=Nx(t),$
with $Q,~{}M\succ 0$. Assuming $(A,B_{u})$ is stabilizable, then we know the solution to this problem is
$N^{\star}=-M^{-1}B_{u}^{\mathrm{T}}Z,$ (30)
where $Z\succ 0$ is the unique positive definite solution of the ARE
$A^{\mathrm{T}}Z+ZA+Q-ZB_{u}R^{-1}B_{u}^{\mathrm{T}}Z=0.$ (31)
Accordingly, a dual setup in deceiving such a defender would be to inject an adversarial perturbation $a(t)=Lx(t)$ in the system, forcing the defender to converge to the incorrect linear-quadratic gain
$N_{a}(L)=-M^{-1}B_{u}^{\mathrm{T}}Z_{a}(L),$ (32)
where $Z_{a}(L)\succ 0$ is obtained by solving the perturbed ARE
$(A+B_{a}L)^{\mathrm{T}}Z_{a}+Z_{a}(A+B_{a}L)\\ +Q-Z_{a}B_{u}R^{-1}B_{u}^{\mathrm{T}}Z_{a}=0.$ (33)
Note that (31) is much easier to analyze than (10); as long as $\mathrm{Ran}(B_{a})\subseteq\mathrm{Ran}(B_{u})$, i.e., as long as the attacker does not have more input channels than the defender, then the positive definite solution to (31) always exists (see [29]). Therefore, it is straightforward to verify that the dual deception problem
$\inf_{L\in\mathbb{R}^{m_{a}\times n},Z_{a}\in\mathbb{R}^{n\times n}}\left\lVert N_{a}(L)-\bar{N}\right\rVert_{\mathrm{F}}^{2}+\mathrm{tr}(L^{\mathrm{T}}\Gamma L),$
[…]
$\qquad\qquad\quad~{}Z_{a}\succ 0,$
admits a global minimizer for any choice of $\Gamma\succ 0$ and $\bar{N}\in\mathbb{R}^{m_{u}\times n}$. With minor modifications, one may then adjust Algorithm 1 to solve (34) numerically (see [29]). The guarantees of convergence will follow the same line as those of Section 5, but will not be constrained by the condition of Lemma 1 since (34) is well-defined for any $L\in\mathbb{R}^{m_{a}\times n}$.
7 Simulations
We consider a linearized model of the ADMIRE benchmark aircraft [32], whose open-loop state and input matrices $A_{o},B_{u}$ are given in the Appendix. Using these, we define the closed-loop state matrix of the aircraft as $A=A_{o}+B_{u}F$, where the gain $F$ originates from a linear-quadratic regulation problem with identity weighting matrices.
7.1 Case 1: Deception against a Performance-Degrading Attack
In the first case, we consider that an adversary has access to the first and the third actuator of the aircraft. Since each actuator corresponds to a column of $B_{u}$, this means that
$\small B_{a}^{\textrm{T}}=\begin{bmatrix}-0.0062&-0.0072&1.2456&2.7172&-0.7497\\ -0.0709&0.0039&-10.6058&-2.4724&-0.4923\\ \end{bmatrix}.$
The adversary then wants to use these actuators to disturb the stabilization of the aircraft. To that end, it seeks to learn the attack that solves the optimization problem (4) with $Q=I_{5}$ and $R=0.6I_{2}$.
Figure 2: Case 1: The evolution of the entries of $\Lambda^{i}$ during Algorithm 1.
Figure 3: Case 1: The evolution of the cost (11) during Algorithm 1.
On the other side, the defender who is aware of the adversary’s objective wants to deceive them to converge as closely as possible to the attack gain $\bar{K}=0$. To that end, the defender plugs a deceptive gain in the system to corrupt the attacker’s learning data. This gain is obtained by performing Algorithm $1$ with $\omega=0.001$, so as to numerically solve the deception optimization (11) with $\Gamma=10^{-3}I_{7}$.
The evolutions of the deception gain sequence $\Lambda^{i}$ generated by Algorithm 1, as well as of the deception cost (11), are shown in Figures 2-3. From Figure 2, we clearly notice that Algorithm 1 reaches a stationary point of the deception problem (11) after about 3500 iterations, which validates Theorem 2. In addition, Figure 3 shows that the deception cost (11) is decreasing at each iteration of Algorithm 1. This is in line with the observation made in Lemma 2, regarding the Algorithm’s equivalence to a first-order gradient method. It indicates that, at each iteration $i$, the gain $\Lambda^{i}$ the defender computes becomes more and more efficient in deceiving the adversary. Table 1 additionally shows the ratio between the optimal attack $K^{\star}$ the attacker would have learned in the absence of deception, and the attack $K_{u}(\Lambda^{\star})$ they were forced to learn instead. Clearly, the latter has been suppressed substantially and is thus more benign in disturbing closed-loop performance. Note that $K_{u}(\Lambda^{\star})$ does not become exactly zero because of the regularization term that penalizes large values of $\Lambda$ in (11).
Table 1: Ratio $[K_{u}(\Lambda^{\star})]_{ij}/[K^{\star}]_{ij}$ between optimal and deceived attack gains.
Col. $1$ Col. $2$ Col. $3$ Col. $4$ Col. $5$
Row $1$ $0.0827$ $0.9429$ $0.0545$ $0.2748$ $0.2439$
Row $2$ $0.1581$ $0.0046$ $0.2179$ $0.1103$ $0.2189$
Table 2: Ratio $[\hat{K}_{u}(\Lambda^{\star})]_{ij}/[\hat{K}^{\star}]_{ij}$ between optimal and deceived attack gains, under incorrectly assumed weighing matrices for the adversary.
Col. $1$ Col. $2$ Col. $3$ Col. $4$ Col. $5$
Row $1$ $0.0655$ $-0.0829$ $0.0355$ $0.2159$ $0.2531$
Row $2$ $0.1190$ $0.0014$ $0.1761$ $0.0694$ $0.1366$
Next, we showcase the robustness of the derived deception gain when the adversary’s weighing matrices are different than those assumed. Specifically, we take the deception gain we previously derived under $Q=I_{5}$ and $R=0.6I_{2}$, and show that even if the adversary’s weighing matrices are equal to $\hat{Q}=5I_{6}\neq Q$ and $\hat{R}=2I_{2}\neq R$, then the learned adversarial attack is still suppressed. We show the ratio between the learned adversarial attack $\hat{K}_{u}(\Lambda^{\star})$ and the nominal attack $\hat{K}^{\star}$ under $\hat{Q},~{}\hat{R}$ in Table 2, validating this conjecture along with the analysis we carried out in Section 6.2.
7.2 Case 2: Deception against a Destabilizing Attack
In Case 2, we consider the same setup for the adversary, but where $R=0.1I_{2}$ in their optimization problem (4). In this case, the optimal attack $K^{\star}$ is destabilizing. To find a deception gain that will force the adversary to learn a more benign attack instead, we re-employ the same scheme as in Case 1 for the defender. However, following the discussion in Section 6, we select $\Lambda^{0}$ in Algorithm $1$ so that it makes the eigenvalues of $A+B_{u}\Lambda^{0}$ sufficiently negative; in particular, equal to $-100$.
The evolution of the deception gain sequence $\Lambda^{i}$ and the deception cost (11) during Algorithm 1 is shown in Figures 4-5. While we still observe a decreasing deception cost and a convergent deception gain, the number of iterations needed to achieve the same level of tolerance increased significantly. This was due to the conservative initialization of $\Lambda^{0}$ very far from the optimal deception gain $\Lambda^{\star}$. Meanwhile, the eigenvalues of the state matrix in closed-loop with the deceptive attack gain are equal to $-17.6911,-7.7253,-2.5628+0.4642i,-2.5628-0.4642i,-1.6568$. This means that the learned optimal attack under deception is not destabilizing, as opposed to the nominal optimal attack $K^{\star}$.
Figure 4: Case 2: The evolution of the entries of $\Lambda^{i}$ during Algorithm 1.
Figure 5: Case 2: The evolution of the cost (11) during Algorithm 1.
8 Conclusion
We study the problem of deception in the context of data-driven linear-quadratic control. Given an adversary (defender) who wants to learn the optimal linear-quadratic attack (control) for a system using data, we design a deceptive input that optimally corrupts this data and forces the learner to converge to a prespecified suboptimal policy. We characterize the deceptive input as the solution of coupled matrix equations, which we subsequently solve using fixed-point iteration and with convergence guarantees.
Future work includes an extension of the presented setup to the deception of nonlinear data-driven optimal control. While such a problem is significantly more challenging, both computationally and mathematically, one can tackle it using partial differential equations to characterize the optimal control, in lieu of a Riccati equation.

### v2

@@
 Despite the heterogeneity of these applications, a common theme among them is that of information asymmetry; that is, to be able to deceive the opponent, one must have an information advantage against them, and the opponent should usually be oblivious to this advantage.
+From a control-theoretic perspective, deception in dynamical systems can be formulated within a broader optimal control and game-theoretic framework.
+In such settings, one agent strategically perturbs information or system inputs in order to influence the policy computed or learned by one or more other agents [6].
+This leads naturally to bilevel optimal control problems in which one decision maker anticipates and shapes the optimization or learning process of another.
 In this paper, we consider deception in the context of data-driven linear-quadratic control.
@@
 A key characteristic of such feedback is that it has the same effect on the attack that the adversary will ultimately learn, irrespective of the adversary’s actual learning algorithm.
-To obtain the defender’s deceptive input, we cast its objective as a constrained optimization problem balancing two objectives: i) the desire to steer the adversary’s learned policy towards an a priori chosen benign gain; and ii) the desire to use as small of deceptive feedback as possible to maintain the nominal closed-loop performance and stability.
+To obtain the defender’s deceptive input, we cast its objective as a constrained optimization problem balancing two objectives: i) the desire to steer the adversary’s learned policy towards an a priori chosen benign gain; and ii) the desire to use as small of deceptive feedback as possible to maintain the nominal closed-loop characteristics and stability.
 We mathematically specify the conditions under which the defender’s optimization is well-defined and characterize its solution as the root of a coupled algebraic Riccati equation and a Lyapunov equation.
@@
 This is a much more challenging problem owing to the associated value function possibly becoming unbounded and thus undefined [27].
+The contributions of the paper are summarized as follows.
+  1. 1.
+    We consider deception in a setting where a defender, who has knowledge of the system’s matrices, injects an additional state-feedback input to strategically influence the learning dynamics of a data-driven attacker, with the goal of steering it toward a suboptimal attack policy.
+Unlike classical LQR game formulations [27, 28, 29], where all players have model access and the equilibrium is uniquely determined by the problem parameters, our setting exploits the fact that the adversary must learn its policy from data, opening the door to shaping what it converges to.
+This also distinguishes our approach from moving target defense [24, 25, 26], which does not explicitly target the adversary’s learned policy, and from deception in Nash equilibrium seeking [11], which considers static game settings.
+  2. 2.
+    We characterize the optimal deception gain as the solution of a coupled algebraic Riccati and Lyapunov equation, and develop a block successive over-relaxation algorithm with provable convergence guarantees.
+This provides theoretical foundations for this specific problem that are absent from related data poisoning approaches, such as those on virtual reference feedback controllers [17, 18] and data predictive control [19], as well as from the preliminary version of this paper [30].
+  3. 3.
+    We analyze robustness to uncertainty in the adversary’s cost matrices and validate the framework on an aircraft model, demonstrating substantial attack suppression.
+Structure: The remainder of the paper is organized as follows.
+Section 2 presents preliminaries, Section 3 formulates the deception problem, and Section 4 outlines the proposed solution within an algorithmic framework.
+Section 5 establishes convergence guarantees for the algorithm, while Section 6 discusses robustness to parametric uncertainty and relaxed assumptions.
+Finally, Section 7 presents simulation results, and Section 8 concludes the paper.
 2 Preliminaries
@@
 The vector $e_{i}\in\mathbb{R}^{n}$ denotes a unit vector with a unity located in the $i$-th row.
-For any matrix $X$, $\left\lVert X\right\rVert_{\mathrm{F}}$ denotes its Frobenius norm, whereas $\mathrm{Ran}(X)$ denotes its range.
-In addition, $[X]_{ij}$ denotes the entry in its $i$-th row and $j$-th column.
+For any matrix $X$, $\left\lVert X\right\rVert_{\mathrm{F}}$ denotes its Frobenius norm, whereas $\mathrm{Ran}(X)$ denotes its range (i.e., its column space).
+In addition, $X]_{ij}$ denotes the entry in its $i$-th row and $j$-th column.
 Moreover, we follow the matrix differentiation convention $\left[\frac{\mathrm{d}}{\mathrm{d}X}\right]_{ij}=\frac{\mathrm{d}}{\mathrm{d}[X]_{ij}}$.
@@
 We use $I_{n}$ to denote the identity matrix of order $n$ (we omit the subscript when the order is obvious).
-Finally, we denote the Kronecker delta function as $\delta_{i,j}$, so that $\delta_{i,j}=1$ if $i=j$, and $\delta_{i,j}=0$ else.
+We denote the Kronecker delta function as $\delta_{i,j}$, so that $\delta_{i,j}=1$ if $i=j$, and $\delta_{i,j}=0$ else.
 Definition 1.
-Consider the following algebraic Riccati equation:
+Consider the algebraic Riccati equation:
 $A^{\mathrm{T}}P+PA+Q-P(B_{1}R_{1}^{-1}B_{1}^{\mathrm{T}}-B_{2}R_{2}^{-1}B_{2}^{\mathrm{T}})P=0,$ (1)
@@
 $a(t)=Kx(t),$
-where $Q,R\succ 0$ are weighting matrices and $K$ is the adversary’s attack gain that is to be optimized.
+where $Q,~R\succ 0$ are weighting matrices and $K$ is the adversary’s attack gain that is to be optimized.
 Provided that $\lambda_{\textrm{min}}(R)$ is above a threshold, it then follows from standard linear systems theory [27] that the solution to this problem is given by
@@
   1. 1.
-    The adversary does not know the values of the entries of the system’s matrices $A,B_{a}$.
+    The adversary does not know the values of the entries of the system matrix $A$. They may or may not know the values of the entries of the system matrix $B_{a}$.
   2. 2.
@@
 We assume that the defender of the system is aware that a potential adversary might be gathering information from the system with the aim of synthesizing the optimal attack gain $K^{\star}$ in (5).
-In that respect, the defender opts to spoof the system’s dynamics during this window, and deceive the adversary into learning a prespecified gain different from $K^{\star}$.
+In response, the defender may temporarily modify the system’s dynamics to induce the adversary to learn a prespecified gain different from $K^{\star}$.
 Specifically, during this window, the system’s dynamics take the form
@@
 where $\bar{K}$ is as in Problem 1, $\Gamma\succ 0$, and $\mathrm{tr}(\Lambda^{\mathrm{T}}\Gamma\Lambda)$ is a regulation term.
-This regulation term is necessary to ensure that the minimum exists, but also to penalize how large the deception gain $\Lambda$ can be to minimize performance loss and maintain stability [^1] This is for the case where $\bar{K}$ is not stabilizing.
+This regulation term is necessary to ensure that the minimum exists, but also to penalize how large the deception gain $\Lambda$ can be to minimize deviation from nominal closed-loop performance and maintain stability [^1] This is for the case where $\bar{K}$ is not stabilizing.
 For the case where $\bar{K}=0$, regularization is not needed; see Section 6.1.
@@
 Input: Tolerance $\epsilon>0$, step size $\omega\in(0,2)$, regulation weight $\Gamma\succ 0$, desired gain $\bar{K}$, parameters $A,B_{u},B_{a},R,Q$.
-      Output: Estimated solution $(\hat{\Lambda},\hat{P}_{u},\hat{\Pi})$ to (14a)-(14c).
+[…]
 1:procedure
@@
 The practical interpretation of Lemma 3 is that, by continuity, $P_{u}(\Lambda)\in\mathbb{S}_{+}$ exists not only everywhere in $\mathcal{S}_{0}$ but also in a neighborhood around it.
-Therefore, the negated gradient $-\frac{\mathrm{d}\tilde{J}(\Lambda)}{\mathrm{d}\Lambda}$ points in a direction towards which $P_{u}(\Lambda)\in\mathbb{S}_{+}$ exists, and hence this direction is in the interior of $\mathcal{S}_{0}$.
+Therefore, the negated gradient $\frac{\mathrm{d}\tilde{J}(\Lambda)}{\mathrm{d}\Lambda}$ points in a direction towards which $P_{u}(\Lambda)\in\mathbb{S}_{+}$ exists, and hence this direction is in the interior of $\mathcal{S}_{0}$.
 Accordingly, Algorithm 1 will always remain within a set where both equations (15) and (16) admit unique solutions and will not terminate prematurely.
@@
 Example 1.
-Consider a scalar system where the closed-loop state matrix is $A=-1$, the input matrices are $B_{u}=B_{a}=1$, and the state cost matrix is $Q=1$.
+Consider a scalar system where the closed-loop state matrix is A=1A=-1, the input matrices are $B_{u}=B_{a}=1$, and the state cost matrix is $Q=1$.
 Then, then ARE (6) reduces to the scalar quadratic equation
@@
 Therefore, from (5), the optimal attack gain is $K^{\star}=1-\sqrt{1-R^{-1}}$.
-On the other hand, when $10<R<$ 1, (26) admits no solutions, the value of (4) is unbounded, and the optimal attack $K^{\star}$ destabilizing.
+On the other hand, when $0<R<1$, (26) admits no solutions, the value of (4) is unbounded, and the optimal attack $K^{\star}$ destabilizing.
 Yet, the defender can deceive the attacker into a milder attack that renders the closed-loop stable, while bypassing the rationale of Lemma 1.
@@
 Then, for every $\epsilon>0$ there exists $\gamma^{\star}>0$ such that if $\Gamma\preceq\gamma^{\star}I$ then $\hat{J}(\Lambda^{\star})\leq\tilde{J}(\Lambda^{\star})+\epsilon$.
+Remark 6.
+Proposition 2 establishes that $\hat{J}(\Lambda^{\star})\leq\tilde{J}(\Lambda^{\star})+\epsilon$.
+The gap $\hat{J}(\Lambda^{\star})-\tilde{J}(\Lambda^{\star})$ depends continuously on the cost-matrix mismatches $\|\hat{Q}-Q\|_{\textrm{F}}$ and $\|\hat{R}-R\|_{\textrm{F}}$ and vanishes as these mismatches vanish.
+This follows from standard norm inequalities together with the continuous dependence of the stabilizing algebraic Riccati solution (and hence of the induced feedback gain) on its cost matrices.
 6.3 On the Deception against Minimizing Data-Driven Linear-Quadratic Regulators
@@
 7 Simulations
-We consider a linearized model of the ADMIRE benchmark aircraft [32], whose open-loop state and input matrices $A_{o},B_{u}$ are given in the Appendix.
+We consider a linearized model of the ADMIRE benchmark aircraft, whose open-loop state and input matrices $A_{o},B_{u}$ are given in [33].
 Using these, we define the closed-loop state matrix of the aircraft as $A=A_{o}+B_{u}F$, where the gain $F$ originates from a linear-quadratic regulation problem with identity weighting matrices.
+Figure 2: Case 1: The evolution of the entries of $\Lambda^{i}$ during Algorithm 1.
+Figure 3: Case 1: The evolution of the cost (11) during Algorithm 1.
+Figure 4: Impact of deception on the evolution of the learning algorithms presented in [34, 35].
+The figure shows the distance of the learnt gain $K_{j}$ at each iteration $j$ of the learning algorithm, from the suboptimal gain $K_{u}(\Lambda^{\star})$.
 7.1 Case 1: Deception against a Performance-Degrading Attack
@@
 To that end, it seeks to learn the attack that solves the optimization problem (4) with $Q=I_{5}$ and $R=0.6I_{2}$.
-Figure 2: Case 1: The evolution of the entries of $\Lambda^{i}$ during Algorithm 1.
-Figure 3: Case 1: The evolution of the cost (11) during Algorithm 1.
 On the other side, the defender who is aware of the adversary’s objective wants to deceive them to converge as closely as possible to the attack gain $\bar{K}=0$.
@@
 The evolutions of the deception gain sequence $\Lambda^{i}$ generated by Algorithm 1, as well as of the deception cost (11), are shown in Figures 2-3.
-From Figure 2, we clearly notice that Algorithm 1 reaches a stationary point of the deception problem (11) after about 3500 iterations, which validates Theorem 2.
+From Figure 2, we notice that Algorithm 1 reaches a stationary point of the deception problem (11) after about 3500 iterations, which validates Theorem 2.
 In addition, Figure 3 shows that the deception cost (11) is decreasing at each iteration of Algorithm 1.
@@
 It indicates that, at each iteration $i$, the gain $\Lambda^{i}$ the defender computes becomes more and more efficient in deceiving the adversary.
+Table 1: Ratio $K_{u}(\Lambda^{\star})]_{ij}/[K^{\star}]_{ij}$ between optimal and deceived attack gains.
+Col. 11 Col. 22 Col. 33 Col. 44 Col. 55
+Row 11 $0.0827$ $0.9429$ $0.0545$ $0.2748$ $0.2439$
+Row 22 $0.1581$ $0.0046$ $0.2179$ $0.1103$ $0.2189$
+Table 2: Ratio $\hat{K}_{u}(\Lambda^{\star})]_{ij}/[\hat{K}^{\star}]_{ij}$ between optimal and deceived attack gains, under incorrectly assumed weighing matrices for the adversary.
+Col. 11 Col. 22 Col. 33 Col. 44 Col. 55
+Row 11 $0.0655$ 0.0829-0.0829 $0.0355$ $0.2159$ $0.2531$
+Row 22 $0.1190$ $0.0014$ $0.1761$ $0.0694$ $0.1366$
+Table 3: State energy $\mathcal{E}_{x}=\int_{0}^{\infty}\left\lVert x(t)\right\rVert^{2}\,dt$ under nominal and deceived attacks for different adversarial input weights $R$.
+The value $\infty$ indicates closed-loop instability.
+Case $R=0.2I_{2}$ $R=0.4I_{2}$ $R=0.6I_{2}$ $R=0.8I_{2}$
+Nominal $\infty$ $0.022844$ $0.008229$ $0.008055$
+Deceived $0.007948$ $0.007924$ $0.007914$ $0.007909$
 Table 1 additionally shows the ratio between the optimal attack $K^{\star}$ the attacker would have learned in the absence of deception, and the attack $K_{u}(\Lambda^{\star})$ they were forced to learn instead.
@@
 Note that $K_{u}(\Lambda^{\star})$ does not become exactly zero because of the regularization term that penalizes large values of $\Lambda$ in (11).
-Table 1: Ratio $[K_{u}(\Lambda^{\star})]_{ij}/[K^{\star}]_{ij}$ between optimal and deceived attack gains.
-Col. $1$ Col. $2$ Col. $3$ Col. $4$ Col. $5$
-Row $1$ $0.0827$ $0.9429$ $0.0545$ $0.2748$ $0.2439$
-Row $2$ $0.1581$ $0.0046$ $0.2179$ $0.1103$ $0.2189$
-Table 2: Ratio $[\hat{K}_{u}(\Lambda^{\star})]_{ij}/[\hat{K}^{\star}]_{ij}$ between optimal and deceived attack gains, under incorrectly assumed weighing matrices for the adversary.
-Col. $1$ Col. $2$ Col. $3$ Col. $4$ Col. $5$
-Row $1$ $0.0655$ $-0.0829$ $0.0355$ $0.2159$ $0.2531$
-Row $2$ $0.1190$ $0.0014$ $0.1761$ $0.0694$ $0.1366$
-Next, we showcase the robustness of the derived deception gain when the adversary’s weighing matrices are different than those assumed.
+To validate this prediction at the level of the attacker’s learning dynamics, Figure 4 illustrates how the deception gain $\Lambda^{\star}$ affects the evolution of two learning algorithms from the literature [34, 35].
+One assumes unknown $A$ and $B_{a}$ matrices [35], whereas the other assumes unknown $A$ and known $B_{a}$ [34].
+In both cases, we observe the learning process converges to the suboptimal attack $K_{u}(\Lambda^{\star})$ that the defender had predicted.
+To further assess the impact at the level of closed-loop performance, Table 3 reports the resulting state energy $\mathcal{E}_{x}=\int_{0}^{\infty}\left\lVert x(t)\right\rVert^{2}dt$ under nominal and deceived attacks for different values of $R$.
+We observe that, for small values of $R$, the nominal attack may even destabilize the system (see next subsection), whereas the deceived attack consistently yields finite and smaller energy.
+Finally, we showcase the robustness of the derived deception gain when the adversary’s weighing matrices are different than those assumed.
 Specifically, we take the deception gain we previously derived under $Q=I_{5}$ and $R=0.6I_{2}$, and show that even if the adversary’s weighing matrices are equal to $\hat{Q}=5I_{6}\neq Q$ and $\hat{R}=2I_{2}\neq R$, then the learned adversarial attack is still suppressed.
 We show the ratio between the learned adversarial attack $\hat{K}_{u}(\Lambda^{\star})$ and the nominal attack $\hat{K}^{\star}$ under $\hat{Q},~\hat{R}$ in Table 2, validating this conjecture along with the analysis we carried out in Section 6.2.
+Figure 5: Case 2: The evolution of the entries of $\Lambda^{i}$ during Algorithm 1.
+Figure 6: Case 2: The evolution of the cost (11) during Algorithm 1.
 7.2 Case 2: Deception against a Destabilizing Attack
@@
 To find a deception gain that will force the adversary to learn a more benign attack instead, we re-employ the same scheme as in Case 1 for the defender.
-However, following the discussion in Section 6, we select $\Lambda^{0}$ in Algorithm $1$ so that it makes the eigenvalues of $A+B_{u}\Lambda^{0}$ sufficiently negative; in particular, equal to $-100$.
-The evolution of the deception gain sequence $\Lambda^{i}$ and the deception cost (11) during Algorithm 1 is shown in Figures 4-5.
+However, following the discussion in Section 6, we select $\Lambda^{0}$ in Algorithm $1$ so that it makes the eigenvalues of $A+B_{u}\Lambda^{0}$ sufficiently negative; in particular, equal to 100-100.
+The evolution of the deception gain sequence $\Lambda^{i}$ and the deception cost (11) during Algorithm 1 is shown in Figures 5-6.
 While we still observe a decreasing deception cost and a convergent deception gain, the number of iterations needed to achieve the same level of tolerance increased significantly.
 This was due to the conservative initialization of $\Lambda^{0}$ very far from the optimal deception gain $\Lambda^{\star}$.
-Meanwhile, the eigenvalues of the state matrix in closed-loop with the deceptive attack gain are equal to $-17.6911,-7.7253,-2.5628+0.4642i,-2.5628-0.4642i,-1.6568$.
+Meanwhile, the eigenvalues of the state matrix in closed-loop with the deceptive attack gain are equal to 17.6911,7.7253,2.5628+0.4642i,2.56280.4642i,1.6568-17.6911,-7.7253,-2.5628+0.4642i,-2.5628-0.4642i,-1.6568.
 This means that the learned optimal attack under deception is not destabilizing, as opposed to the nominal optimal attack $K^{\star}$.
-Figure 4: Case 2: The evolution of the entries of $\Lambda^{i}$ during Algorithm 1.
-Figure 5: Case 2: The evolution of the cost (11) during Algorithm 1.
 8 Conclusion